What couple must be applied to a 1 m long wire with 1-mm diameter in order to twist one end of it through 90°, the other end remaining fixed? The rigidity modulus is 2.8 x 10¹⁰ N/m^2


1
Expert's answer
2020-12-17T07:26:32-0500
"C=\frac{\pi N r^4}{2l}\theta\\\theta=0.5\pi\\C=\frac{\pi^2(2.8\cdot10^{10})(5\cdot10^{-4})^4}{4(1)}=4.3\cdot10^{-3}\ Nm"